OpenStudy (anonymous):

What is the limit as x is approaching infinity of (ln(ln(x)))/x Please explain how you got it too!

OpenStudy (agreene):

Use L'Hopitals and get: \[\lim_{x \rightarrow \infty}\frac{1}{x\ln (x)}\] the denominator goes to infty and \[\frac{1}{\infty}\rightarrow0\]